A shipment of 13 microwave ovens contains 4 defective units. A restaurant buys three of these units. What is the probability of the restaurant by at least two non defective units? The probability of the restaurant buying at least two nondefective units is
Answer #1

P(Non Defective) = 9/13 = 0.6923

This is a binomial distribution question with

n = 3

p = 0.6923

q = 1 - p = 0.3077

where

Please hit thumps up if the answer helped you